Frame Masters, Inc., recently purchased $5500 worth of new equipment in order to offer a new style of eyeglass frame. The marketing department of Frame Masters estimates that the cost of producing this new frame is $18 and that the frame will be sold to stores for $38. Find the number of frames that must be sold in order to break even
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)




Let x as the number of frames sold

R(x) the revenue from selling x frames
C(x) as the total cost of producing and selling x frames

Revenue = 38x
Cost = 5000+18x
For break even
38x = 5000+18x
20x = 5000
x= 275
Sell 275 frames to break even.







Answer by ikleyn(52813)

.

Let x be the number of frames sold.

R(x) the revenue from selling x frames.
C(x) as the total cost of producing and selling x frames.

Revenue = 38x
Cost = 5500+18x
For break even
38x = 5500+18x
20x = 5500
x= 275

Sell 275 frames to break even.